Assuming that 80% of the Mars surface is covered with water at an average




depth of 3.5 miles and the radius of the earth is given by 5.97x106 meter,




estimate the mass of the water on the Mars in kilogram

1
Expert's answer
2022-01-23T16:15:30-0500

"m=0.8\rho V=0.8\rho\frac43\pi(R^3-(R-r)^3)=60\cdot 10^{18}~kg."